# Exactly or Approximately Wasserstein Distributionally Robust Estimation According to Wasserstein Radii Being Small or Large

### v1

@@ 1 Introduction
Notation: Let $\mathbb{R}^{n}$ be the $n$ dimensional real vector space and $\mathbb{R}^{n\times m}$ be the $n\times m$ dimensional real matrix space. The notation $I_{n}$ stands for the identity matrix in $\mathbb{R}^{n\times n}$ and $\bf{0}$ denotes the zero matrix of proper dimension. For any $A\in\mathbb{R}^{n\times n}$, we use $A^{-}$ and $A^{\dagger}$ to denote the generalized inverse and the Moore-Penrose pseudo-inverse of the matrix AA. The notation $\mbox{Tr}(A)$ denotes the trace of the matrix AA and the null space of AA is denoted by $\mbox{Null}(A)$. For any $A,B\in\mathbb{R}^{n\times n}$, the inner product of AA and $B$ is denoted by $\left\langle A,B\right\rangle=\mbox{Tr}(AB)$ and $A\perp B$ means $\left\langle A,B\right\rangle=0$. Let $\mathbb{S}^{n}$ be the space of symmetric matrices in $\mathbb{R}^{n\times n}$ and $\mathbb{S}^{n}_{+}$ be the cone of positive semidefinite matrices in $\mathbb{S}^{n}$. For any $A\in\mathbb{S}^{n}$, we use $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ to denote the minimum and maximum eigenvalues of AA, respectively. For any $A\in\mathbb{S}^{n}_{+}$, the notation $A^{\frac{1}{2}}$ denotes the unique positive semidefinite square root of AA. For any $A,B\in\mathbb{S}^{n}$, the relation $A\succ B$ $(A\succeq B)$ means that ABA-B is positive definite (semidefinite). For a measurable function of two variables $f(x,y)$, we use $\bigtriangledown_{x}f(x,y)$ to denote the partial derivative of $f(x,y)$ with respect to $x$. The notation $\|\cdot\|$ denotes the Euclidean norm. A normal distribution with mean vector $\mu$ and covariance matrix $\Sigma$ is denoted by $\mathcal{N}(\mu,\Sigma)$.
@@ 3.1 An Finite-dimensional Problem Related to the Existence of the Saddle Point Solution for (WDRE)
where the objective function is given by (3), and the constraint sets are defined by (6) and (5), respectively. Essentially, (LG-WDRE) is a finite-dimensional robust estimation problem, because the linear estimator $f$ can be parameterized by the matrix AA and the vector $b$, and the Gaussian distributions can also be fully described by their mean vectors and covariance matrices. Next, we shall explore the equivalence between the existence of the saddle point solution in (WDRE) and that in (LG-WDRE).
@@ 3.2 A Necessary and Sufficient Condition for the Existence of the Saddle Point Solution for (WDRE)
For fixed $\mu_{x}$, $\mu_{w}$, $\Sigma_{x}$, $\Sigma_{w}$ and AA, the optimal solution to the inner minimization problem is given by $b^{*}=(I_{n}-AH)\mu_{x}-A\mu_{w}$, which minimizes the objective function by eliminating the quadratic term with respect to $b$. By substituting $b^{*}$ into the above equation, problem (13) becomes
@@
Since the objective function of (14) is convex in AA for each $(\Sigma_{x},\Sigma_{w})$, concave in $(\Sigma_{x},\Sigma_{w})$ for each AA, and $\mathbb{B}_{\Sigma}$ is convex and compact kuhn2019wasserstein , we can apply Sion’s minimax theorem sion1958general to conclude that
@@
where (15) is a reformulation of (14) with simplified notations. Furthermore, since $\mathbb{B}_{\Sigma}$ is compact, and there exists a pair of positive definite matrices $(\Sigma_{x},\Sigma_{w})\in\mathbb{B}_{\Sigma}$ such that the objective function is coercive in AA, then the above problems admit the saddle point solution (botelho2014functional, , Theorem 10.2).
@@
On the other hand, for any fixed AA, consider the inner maximization problem in (16)
@@
Since $(\Sigma_{x}^{*},\Sigma_{w}^{*},A^{*},\gamma_{x}^{*},\gamma_{w}^{*})$ is a saddle point solution of (17), for any $\Sigma_{x}\succeq{\bf 0}$, $\Sigma_{w}\succeq{\bf 0}$, AA, $\gamma_{x}\geq 0$ and $\gamma_{w}\geq 0$, we have
@@
Then for any $\Sigma_{x}\succeq{\bf 0}$, $\Sigma_{w}\succeq{\bf 0}$, AA, $\gamma_{x}\geq 0$ and $\gamma_{w}\geq 0$, the objective function of (17), defined as $G$ in (21), satisfies
@@ 4.1 A Tight Convex Relaxation of (LG-WDRE)
For given AA and $b$, the objective function of (49) is convex in $(\mu_{x},\mu_{w})$. Consequently, the inner maximization problem over $(\mu_{x},\mu_{w})$ is non-convex, which brings a significant challenge in identifying the optimal solution. For ease of notations, we denote $\tilde{\mu}_{x}\triangleq\mu_{x}-\hat{\mu}_{x}$, $\tilde{\mu}_{w}\triangleq\mu_{w}-\hat{\mu}_{w}$ and $\tilde{b}\triangleq b+(AH-I_{n})\hat{\mu}_{x}+A\hat{\mu}_{w}$, respectively. Then problem (49) can be reformulated as follows
@@
Notice that for given AA, $\tilde{b}$, $\Sigma_{x}$ and $\Sigma_{w}$, the inner maximization problem over $(\tilde{\mu}_{x},\tilde{\mu}_{w})$ in (50) is a nonhomogeneous QCQP problem with two homogeneous constraints. If we directly apply the SDP relaxation method, we can not assert that the relaxed problem is equivalent to problem (50) ye2003new ; ai2024tightness . To address this difficulty, we first give a special structure of the optimal solution to (50) in the following theorem.
@@
In addition, for given AA and $\tilde{b}$, the optimal solution to the inner maximization problem in (50) is denoted as $)(\tilde{\mu}_{x}^{*}(A,\tilde{b}),\tilde{\mu}_{w}^{*}(A,\tilde{b}),\Sigma_{x}^{*}(A,\tilde{b}),\Sigma_{w}^{*}(A,\tilde{b}$)). Then for any AA and $\tilde{b}\neq{\bf 0}$, we have
@@
where the first inequality holds because: i) the first two terms of the objective function $\psi$ remain unchanged; ii) $(\tilde{\mu}_{x}^{*}(A,{\bf 0}),\tilde{\mu}_{w}^{*}(A,{\bf 0}),\Sigma_{x}^{*}(A,{\bf 0}),\Sigma_{w}^{*}(A,{\bf 0}))$ is a feasible solution and so is $(-\tilde{\mu}_{x}^{*}(A,{\bf 0}),-\tilde{\mu}_{w}^{*}(A,{\bf 0}),\Sigma_{x}^{*}(A,{\bf 0}),\Sigma_{w}^{*}(A,{\bf 0}))$; iii) at least one of these two feasible solutions satisfies $\left[(AH-I_{n})\tilde{\mu}_{x}+A\tilde{\mu}_{w}\right]^{\top}\tilde{b}\geq 0$; iv) $\tilde{b}^{\top}\tilde{b}>0$, and the second inequality follows from the optimality of $)(\tilde{\mu}_{x}^{*}(A,\tilde{b}),\tilde{\mu}_{w}^{*}(A,\tilde{b}),\Sigma_{x}^{*}(A,\tilde{b}),\Sigma_{w}^{*}(A,\tilde{b}$)) for AA and $\tilde{b}$. Therefore, the optimal solution to (50) must satisfy $\tilde{b}^{*}=\bf 0$. ∎
@@
Notice that for given AA, $\Sigma_{x}$ and $\Sigma_{w}$, the inner maximization problem over $(\tilde{\mu}_{x},\tilde{\mu}_{w})$ in (51) is a homogeneous QCQP problem with two homogeneous constraints, which can be be equivalently reformulated as a convex problem through the SDP relaxation method as described in the following theorem.
@@
Notice that for any AA, $\Sigma_{x}$ and $\Sigma_{w}$, the inner maximization problem on $S$ in (67) always admits a rank-one optimal solution (ye2003new, , Theorem 2.5). This implies that the relaxation is tight, i.e., problems (51) and (67) are equivalent.
Note that the objective function of (67) is convex in AA for each $(S,\Sigma_{x},\Sigma_{w})$, concave in $(S,\Sigma_{x},\Sigma_{w})$ for each AA, and the constraint set under the supremum is convex and compact. Then according to Sion’s minimax theorem sion1958general , (67) is equivalent to the following problem obtained by exchanging minimization and maximization:
@@
Then since the constraints of (72) are independent of AA, we first consider the inner unconstrained minimization problem

### v2

@@ 1 Introduction
Notation: Let $\mathbb{R}^{n}$ be the $n$ dimensional real vector space and $\mathbb{R}^{n\times m}$ be the $n\times m$ dimensional real matrix space. The notation $I_{n}$ stands for the identity matrix in $\mathbb{R}^{n\times n}$ and $\bf{0}$ denotes the zero matrix of proper dimension. For any $A\in\mathbb{R}^{n\times n}$, we use $A^{-}$ and $A^{\dagger}$ to denote the generalized inverse and the Moore-Penrose pseudo-inverse of the matrix $A$. The notation $\mbox{Tr}(A)$ denotes the trace of the matrix $A$ and the null space of $A$ is denoted by $\mbox{Null}(A)$. For any $A,B\in\mathbb{R}^{n\times n}$, the inner product of $A$ and $B$ is denoted by $\left\langle A,B\right\rangle=\mbox{Tr}(AB)$ and $A\perp B$ means $\left\langle A,B\right\rangle=0$. Let $\mathbb{S}^{n}$ be the space of symmetric matrices in $\mathbb{R}^{n\times n}$ and $\mathbb{S}^{n}_{+}$ be the cone of positive semidefinite matrices in $\mathbb{S}^{n}$. For any $A\in\mathbb{S}^{n}$, we use $\lambda_{\min}(A)$ and $\lambda_{\max}(A)$ to denote the minimum and maximum eigenvalues of $A$, respectively. For any $A\in\mathbb{S}^{n}_{+}$, the notation $A^{\frac{1}{2}}$ denotes the unique positive semidefinite square root of $A$. For any $A,B\in\mathbb{S}^{n}$, the relation $A\succ B$ $(A\succeq B)$ means that ABA-B is positive definite (semidefinite). For a measurable function of two variables $f(x,y)$, we use $\bigtriangledown_{x}f(x,y)$ to denote the partial derivative of $f(x,y)$ with respect to $x$. The notation $\|\cdot\|$ denotes the Euclidean norm. A normal distribution with mean vector $\mu$ and covariance matrix $\Sigma$ is denoted by $\mathcal{N}(\mu,\Sigma)$.
@@ 3.1 An Finite-dimensional Problem Related to the Existence of the Saddle Point Solution for (WDRE)
where the objective function is given by (3), and the constraint sets are defined by (6) and (5), respectively. Essentially, (LG-WDRE) is a finite-dimensional robust estimation problem, because the linear estimator $f$ can be parameterized by the matrix $A$ and the vector $b$, and the Gaussian distributions can also be fully described by their mean vectors and covariance matrices. Next, we shall explore the equivalence between the existence of the saddle point solution in (WDRE) and that in (LG-WDRE).
@@ 3.2 A Necessary and Sufficient Condition for the Existence of the Saddle Point Solution for (WDRE)
For fixed $\mu_{x}$, $\mu_{w}$, $\Sigma_{x}$, $\Sigma_{w}$ and $A$, the optimal solution to the inner minimization problem is given by $b^{*}=(I_{n}-AH)\mu_{x}-A\mu_{w}$, which minimizes the objective function by eliminating the quadratic term with respect to $b$. By substituting $b^{*}$ into the above equation, problem (13) becomes
@@
Since the objective function of (14) is convex in $A$ for each $(\Sigma_{x},\Sigma_{w})$, concave in $(\Sigma_{x},\Sigma_{w})$ for each $A$, and $\mathbb{B}_{\Sigma}$ is convex and compact kuhn2019wasserstein , we can apply Sion’s minimax theorem sion1958general to conclude that
@@
where (15) is a reformulation of (14) with simplified notations. Furthermore, since $\mathbb{B}_{\Sigma}$ is compact, and there exists a pair of positive definite matrices $(\Sigma_{x},\Sigma_{w})\in\mathbb{B}_{\Sigma}$ such that the objective function is coercive in $A$, then the above problems admit the saddle point solution (botelho2014functional, , Theorem 10.2).
@@
On the other hand, for any fixed $A$, consider the inner maximization problem in (16)
@@
Since $(\Sigma_{x}^{*},\Sigma_{w}^{*},A^{*},\gamma_{x}^{*},\gamma_{w}^{*})$ is a saddle point solution of (17), for any $\Sigma_{x}\succeq{\bf 0}$, $\Sigma_{w}\succeq{\bf 0}$, $A$, $\gamma_{x}\geq 0$ and $\gamma_{w}\geq 0$, we have
@@
Then for any $\Sigma_{x}\succeq{\bf 0}$, $\Sigma_{w}\succeq{\bf 0}$, $A$, $\gamma_{x}\geq 0$ and $\gamma_{w}\geq 0$, the objective function of (17), defined as $G$ in (21), satisfies
@@ 4.1 A Tight Convex Relaxation of (LG-WDRE)
For given $A$ and $b$, the objective function of (49) is convex in $(\mu_{x},\mu_{w})$. Consequently, the inner maximization problem over $(\mu_{x},\mu_{w})$ is non-convex, which brings a significant challenge in identifying the optimal solution. For ease of notations, we denote $\tilde{\mu}_{x}\triangleq\mu_{x}-\hat{\mu}_{x}$, $\tilde{\mu}_{w}\triangleq\mu_{w}-\hat{\mu}_{w}$ and $\tilde{b}\triangleq b+(AH-I_{n})\hat{\mu}_{x}+A\hat{\mu}_{w}$, respectively. Then problem (49) can be reformulated as follows
@@
Notice that for given $A$, $\tilde{b}$, $\Sigma_{x}$ and $\Sigma_{w}$, the inner maximization problem over $(\tilde{\mu}_{x},\tilde{\mu}_{w})$ in (50) is a nonhomogeneous QCQP problem with two homogeneous constraints. If we directly apply the SDP relaxation method, we can not assert that the relaxed problem is equivalent to problem (50) ye2003new ; ai2024tightness . To address this difficulty, we first give a special structure of the optimal solution to (50) in the following theorem.
@@
In addition, for given $A$ and $\tilde{b}$, the optimal solution to the inner maximization problem in (50) is denoted as $)(\tilde{\mu}_{x}^{*}(A,\tilde{b}),\tilde{\mu}_{w}^{*}(A,\tilde{b}),\Sigma_{x}^{*}(A,\tilde{b}),\Sigma_{w}^{*}(A,\tilde{b}$)). Then for any $A$ and $\tilde{b}\neq{\bf 0}$, we have
@@
where the first inequality holds because: i) the first two terms of the objective function $\psi$ remain unchanged; ii) $(\tilde{\mu}_{x}^{*}(A,{\bf 0}),\tilde{\mu}_{w}^{*}(A,{\bf 0}),\Sigma_{x}^{*}(A,{\bf 0}),\Sigma_{w}^{*}(A,{\bf 0}))$ is a feasible solution and so is $(-\tilde{\mu}_{x}^{*}(A,{\bf 0}),-\tilde{\mu}_{w}^{*}(A,{\bf 0}),\Sigma_{x}^{*}(A,{\bf 0}),\Sigma_{w}^{*}(A,{\bf 0}))$; iii) at least one of these two feasible solutions satisfies $\left[(AH-I_{n})\tilde{\mu}_{x}+A\tilde{\mu}_{w}\right]^{\top}\tilde{b}\geq 0$; iv) $\tilde{b}^{\top}\tilde{b}>0$, and the second inequality follows from the optimality of $)(\tilde{\mu}_{x}^{*}(A,\tilde{b}),\tilde{\mu}_{w}^{*}(A,\tilde{b}),\Sigma_{x}^{*}(A,\tilde{b}),\Sigma_{w}^{*}(A,\tilde{b}$)) for $A$ and $\tilde{b}$. Therefore, the optimal solution to (50) must satisfy $\tilde{b}^{*}=\bf 0$. ∎
@@
Notice that for given $A$, $\Sigma_{x}$ and $\Sigma_{w}$, the inner maximization problem over $(\tilde{\mu}_{x},\tilde{\mu}_{w})$ in (51) is a homogeneous QCQP problem with two homogeneous constraints, which can be be equivalently reformulated as a convex problem through the SDP relaxation method as described in the following theorem.
@@
Notice that for any $A$, $\Sigma_{x}$ and $\Sigma_{w}$, the inner maximization problem on $S$ in (67) always admits a rank-one optimal solution (ye2003new, , Theorem 2.5). This implies that the relaxation is tight, i.e., problems (51) and (67) are equivalent.
Note that the objective function of (67) is convex in $A$ for each $(S,\Sigma_{x},\Sigma_{w})$, concave in $(S,\Sigma_{x},\Sigma_{w})$ for each $A$, and the constraint set under the supremum is convex and compact. Then according to Sion’s minimax theorem sion1958general , (67) is equivalent to the following problem obtained by exchanging minimization and maximization:
@@
Then since the constraints of (72) are independent of $A$, we first consider the inner unconstrained minimization problem